Solve the following quadratic equation using the formula method:
7x/3x-4+2x/6-x=10
1
Expert's answer
2019-06-05T05:33:57-0400

We reduce fractions to a common denominator and get


"\frac {7x} {3x-4} +\frac {2x} {6-x} - 10=0"

"\frac {7x(6-x)} {(3x-4)(6-x)} +\frac {2x(3x-4)} {(6-x)(3x-4)} - \frac {10(6-x)(3x-4)} {(6-x)(3x-4)}=0"


"\frac {42x-7x^2+6x^2-8x+30x^2+180x-240+40x}{(6-x)(3x-4)}=0"

"\frac {29x^2 - 186x + 240}{(6-x)(3x-4)}=0"

A fraction is zero if its numerator is zero.


"29x^2 - 186x + 240=0"

Solve the quadratic equation. Find the discriminant


"D=186^2-4\times 29 \times 240=34596-27840=6756"

Therefore the roots of the equation are equal


"x_1=\frac {186-\sqrt {6756}} {2\times 29}= \frac {93-\sqrt {1689}} { 29}"

"x_2=\frac {186+\sqrt {6756}} {2\times 29}= \frac {93+\sqrt {1689}} { 29}"

Answer.

"x_1= \frac {93-\sqrt {1689}} { 29}"

"x_2= \frac {93+\sqrt {1689}} { 29}"